OpenStudy (anonymous):

Find the ratio of m to n. 9n=3m

OpenStudy (anonymous):

divide both sides by n 9n/n=3m/n ----------> 9=3m/n divide both sides by 3 9/3=3m/3n ---------> 3=m/n so the ratio of m to n is 3:1